A pool company has learned​ that, by pricing a newly released noodle at ​$3​,
sales will reach 15,000 noodles per day during the summer. Raising the price to ​$5 will cause the sales to fall to 13,000 noodles per day.​ [Hint: The line must pass through ​(3,15000​) and ​(5,13000).]
a. Assume that the relationship between sales price and number of noodles sold is linear and write an equation describing this relationship.
B. Predict the daily sales of noodles if the price is ​$3.50.

Answer by Boreal(15235)

slope is (13000-15000)/(5-3)=-2000/2=-1000
point slope formula: y-y1=m(x-x1), m slope and (x1,y1) a point
y-15000=-1000(x-3)
y-15000=-1000x+3000
y=-1000x+18000
when x=$3.50
y=-3500+18000=14,500 noodles predicted to be sold.